# A rigidity theorem for complex Kleinian groups

### v1

@@ 4.2. Ruling out compact simple factors
Using the previous lemma, we provide a sufficient condition for the real Zariski closure of an irreducible $\mathsf{P}_{k}$-Anosov subgroup to have no compact simple factors.
Lemma 4.5.
Suppose $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a strongly irreducible $\mathsf{P}_{k}$-Anosov subgroup and $\mathsf{G}$ is the real Zariski closure of $\Gamma$. If either
  • $k=1$ or
  • $k>1$ and $\Gamma$ is also $\mathsf{P}_{k-1}$-Anosov,
then $\operatorname{\mathsf{G}}$ has no compact simple factors.
Proof.
It suffices to prove the lemma for the identity component $\operatorname{\mathsf{G}}^{0}$ of $\operatorname{\mathsf{G}}$. Since $\Gamma$ is strongly irreducible, $\Gamma\cap\operatorname{\mathsf{G}}^{0}$ is irreducible and its real Zariski closure is $\operatorname{\mathsf{G}}^{0}$. By part (1) of Lemma 4.3, $\operatorname{\mathsf{G}}^{0}$ has trivial center, so we may decompose $\operatorname{\mathsf{G}}^{0}$ into a direct product of simple Lie groups. In particular, we can write
$\operatorname{\mathsf{G}}^{0}=\mathsf{H}\times\mathsf{L}$
where $\mathsf{H}$ is the product of the non-compact simple factors and $\mathsf{L}$ is the product of the compact simple factors. Let $\mathsf{H}^{\prime}$ and $\mathsf{L}^{\prime}$ respectively denote the preimages of $\mathsf{H}$ and $\mathsf{L}$ under the quotient map $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})\rightarrow\mathsf{PSL}(d,\operatorname{\mathbb{K}})$.
Suppose that $\mathsf{L}$ is not the trivial group. By Lemma 4.4, there exist integers $d_{1}$ and $d_{2}$ larger than $1$, representations $\rho_{1}:\mathsf{H}^{\prime}\rightarrow\operatorname{\mathsf{SL}}(d_{1},\operatorname{\mathbb{K}})$ and $\rho_{2}:\mathsf{L}^{\prime}\rightarrow\operatorname{\mathsf{SL}}(d_{2},\operatorname{\mathbb{K}})$ and an identification
such that an element $g=h\ell\in\mathsf{H}^{\prime}\mathsf{L}^{\prime}$ acts by
$g(v\otimes w)=(\rho_{1}(h)(v))\otimes(\rho_{2}(\ell)(w)).$
By conjugating we can assume that $\rho_{2}(\mathsf{L}^{\prime})\subset\mathsf{PU}(d_{2})$. Then endowing $\operatorname{\mathbb{K}}^{d_{1}}\otimes\operatorname{\mathbb{K}}^{d_{2}}$ with the inner product defined by
$\left\langle v_{1}\otimes w_{1},v_{2}\otimes w_{2}\right\rangle=\left\langle v_{1},v_{2}\right\rangle\left\langle w_{1},w_{2}\right\rangle,$
the associated singular values of an element $h\ell\in\mathsf{H}^{\prime}\mathsf{L}^{\prime}$ are
$\mu_{1}(\rho_{1}(h)),\dots,\mu_{1}(\rho_{1}(h)),\mu_{2}(\rho_{1}(h)),\dots,\mu_{2}(\rho_{1}(h)),\dots,\mu_{d_{1}}(\rho_{1}(h)),\dots,\mu_{d_{1}}(\rho_{1}(h))$
with each singular value repeated $d_{2}$ times. Hence $\Gamma$ cannot be $\operatorname{\mathsf{P}}_{1}$-Anosov or $\operatorname{\mathsf{P}}_{k-1,k}$-Anosov for some $k>1$. So we have a contradiction. ∎
4.3. Ruling out multiple non-compact simple factors
Next, we prove that if $\Gamma$ satisfies the hypothesis Theorem 4.1, then the real Zariski closure of $\Gamma$ cannot have multiple non-compact simple factors. To do so, we will use the following result.
Theorem 4.6.
([9, Thm. 1.5]) Suppose $\mathsf{H}$ is the identity component of a semisimple real algebraic Lie group and that $\Gamma<\mathsf{H}$ is $\mathsf{P}_{\theta}$-Anosov and Zariski dense. If $\Lambda_{\theta}(\Gamma)$ is a ${\mathcal{C}}^{1}$-submanifold of $\mathcal{F}_{\theta}$, then $\mathsf{P}_{\theta}$ is a maximal parabolic subgroup.
To apply Theorem 4.6 we first need to establish a few lemmas. The first is an immediate consequence of a result of Labourie [25, Prop. 10.3].
Lemma 4.7.
If $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a strongly irreducible subgroup and $V_{1},\dots,V_{m},W_{1},\dots,W_{m}\subset\operatorname{\mathbb{K}}^{d}$ are linear subspaces with
$\dim V_{j}+\dim W_{j}\leq d$
for all $j=1,\dots,m$, then there exists $\gamma\in\Gamma$ where
$V_{j}\cap\gamma W_{j}=\{0\}$
for all $j=1,\dots,m$.
Proof.
Let $\mathsf{G}\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ denote the real Zariski closure of $\Gamma$ and let $\operatorname{\mathsf{G}}^{0}$ denote the connected component of the identity with respect to the Zariski topology in $\operatorname{\mathsf{G}}$. Since $\Gamma$ is strongly irreducible, $\operatorname{\mathsf{G}}^{0}$ is irreducible. By [25, Prop. 10.3] (this is stated for $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$, but the same arguments works for $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$) the set
$\mathcal{O}_{j}:=\{g\in\operatorname{\mathsf{G}}^{0}:V_{j}\cap gW_{j}=\{0\}\}$
is non-empty and also, by definition, open in the real Zariski topology. So $\mathcal{O}:=\cap_{j=1}^{m}\mathcal{O}_{j}$ is also non-empty and open in the real Zariski topology. Since $\Gamma\cap\operatorname{\mathsf{G}}^{0}$ is Zariski dense in $\operatorname{\mathsf{G}}^{0}$, then there exists some $\gamma\in\Gamma\cap\mathcal{O}$. ∎
The second guarantees the compactness of the subgroup of $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ that fixes the limit set of a strongly irreducible $\operatorname{\mathsf{P}}_{k}$-Anosov subgroup.
Lemma 4.8.
Suppose $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a strongly irreducible $\mathsf{P}_{k}$-Anosov subgroup and
$\mathsf{S}:=\{g\in\mathsf{PSL}(d,\operatorname{\mathbb{K}}):gx=x\text{ for all }x\in\Lambda_{k}(\Gamma)\}.$
Then
  • If $k=1$, then $\mathsf{S}$ is trivial.
  • If $k>1$, then $\mathsf{S}$ is compact.
Proof.
When $k=1$, $\mathsf{S}$ is trivial, since $\Lambda_{1}(\Gamma)$ contains a projective frame (see [7, Lemma 2.18]) and an element of $\mathsf{PSL}(d,\mathbb{K})$ is determined by its action on a projective frame.
Now assume $k>1$. Suppose for a contradiction that $\mathsf{S}$ is not compact. Then we can find an unbounded sequence $\{s_{n}\}\subset\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})$ with $\{[s_{n}]\}\subset\mathsf{S}$. Let $s_{n}=m_{n}a_{n}\ell_{n}$ be a singular value decomposition, i.e $m_{n},\ell_{n}\in\mathsf{SU}(d,\operatorname{\mathbb{K}})$ and $a_{n}=\exp(\kappa(s_{n}))$. Since $\mathsf{SU}(d,\operatorname{\mathbb{K}})$ is compact, by passing to a subsequence we can suppose that $m_{n}\rightarrow m$ and $\ell_{n}\rightarrow\ell$. We can also assume that
$\frac{\sigma_{i_{1}}(s_{n})\cdots\sigma_{i_{k}}(s_{n})}{\sigma_{1}(s_{n})\cdots\sigma_{k}(s_{n})}\rightarrow a_{i_{1},\dots,i_{k}}\in[0,1]$
for all $1\leq i_{1}<\cdots<i_{k}\leq d$.
For any $i\in\{1,\dots,d\}$, let $\eta_{i}:\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})\to\operatorname{\mathsf{SL}}(\wedge^{i}\operatorname{\mathbb{R}}^{d})$ denote the representation induced by the linear action of $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$ on $\wedge^{k}\operatorname{\mathbb{R}}^{d}$. The standard basis $\{e_{1},\dots,e_{d}\}$ of $\operatorname{\mathbb{K}}^{d}$ induces a standard basis
$\{e_{i_{1}}\wedge\cdots\wedge e_{i_{k}}:1\leq i_{1}<\dots<i_{k}\leq d\}$
of $\wedge^{k}\operatorname{\mathbb{K}}^{d}$. Then let $D_{i_{1},\dots,i_{k}}\in{\rm End}(\wedge^{k}\operatorname{\mathbb{K}}^{d})$ be the linear map that fixes the vector $e_{i_{1}}\wedge\dots\wedge e_{i_{k}}$ and sends all other standard basis vectors to 0. Observe that
$\frac{\eta_{k}(s_{n})}{\sigma_{1}(s_{n})\cdots\sigma_{k}(s_{n})}\rightarrow T:=\eta_{k}(m)\left(\sum_{1\leq i_{1}<\cdots<i_{k}\leq d}a_{i_{1},\dots,i_{k}}D_{i_{1},\dots,i_{k}}\right)\eta_{k}(\ell)$
in ${\rm End}(\wedge^{k}\operatorname{\mathbb{K}}^{d})$. Further, if $x\in\operatorname{\mathbb{P}}(\wedge^{k}\operatorname{\mathbb{K}}^{d})\smallsetminus\operatorname{\mathbb{P}}(\ker T)$, then
$T(x)=\lim_{n\rightarrow\infty}\eta_{k}(s_{n})x.$ (5)
Now we view $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})$ as a subset of $\operatorname{\mathbb{P}}(\wedge^{k}\operatorname{\mathbb{K}}^{d})$ using the Plücker embedding, i.e. we identify $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})$ with the image of the embedding $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})\to\operatorname{\mathbb{P}}(\wedge^{k}\operatorname{\mathbb{K}}^{d})$ given by $x\mapsto[h_{1}\wedge\dots\wedge h_{k}]$, where $\{h_{1},\dots,h_{k}\}$ is some (any) basis of $x$. Since $\mathsf{S}$, and hence $\eta_{k}(\mathsf{S})$, fixes $\Lambda_{k}(\Gamma)$ pointwise, Equation (5) implies that
$x=T(x)\quad\text{for all }\quad x\in\Lambda_{k}(\Gamma)\smallsetminus\operatorname{\mathbb{P}}(\ker T).$ (6)
Also, notice that $a_{1,\dots,k}=1$ and so
$\ker T$ $\subset\eta_{k}(\ell)^{-1}\{v\in\wedge^{k}\operatorname{\mathbb{R}}^{d}:v\wedge e_{k+1}\wedge\cdots\wedge e_{d}=0\}$ (7)
$=\{v\in\wedge^{k}\operatorname{\mathbb{R}}^{d}:v\wedge\eta_{d-k}(\ell)^{-1}(e_{k+1}\wedge\cdots\wedge e_{d})=0\}.$
Fix $x_{0}\in\Lambda_{k}(\Gamma)$, and choose vectors $v_{1},\dots,v_{k}\in\operatorname{\mathbb{K}}^{d}$ such that $x_{0}=[v_{1}\wedge\dots\wedge v_{k}]$. Since $\{s_{n}\}$ is unbounded, there exists $1\leq i_{1}^{\prime}<\cdots<i_{k}^{\prime}\leq d$ with $a_{i_{1}^{\prime},\dots,i_{k}^{\prime}}=0$. Then let $1\leq j_{1}<\cdots<j_{d-k}\leq d$ be the indices with
$\{i_{1}^{\prime},\dots,i_{k}^{\prime},j_{1},\dots,j_{d-k}\}=\{1,\dots,d\}.$
By Lemma 4.7 (applied with $W_{1}=W_{2}=\operatorname{Span}(v_{1},\dots,v_{k})$, $V_{1}=\ell^{-1}\operatorname{Span}(e_{k+1},\dots,e_{d})$, and $V_{2}=m\operatorname{Span}(e_{j_{1}},\dots,e_{j_{d-k}})$), there exists $\gamma\in\Gamma$ such that
$\eta_{k}(\gamma)(v_{1}\wedge\cdots\wedge v_{k})\wedge\eta_{d-k}(\ell)^{-1}(e_{k+1}\wedge\cdots\wedge e_{d})\neq 0$ (8)
and
$\eta_{k}(\gamma)(v_{1}\wedge\cdots\wedge v_{k})\wedge\eta_{d-k}(m)(e_{j_{1}}\wedge\cdots\wedge e_{j_{d-k}})\neq 0.$ (9)
Let $v:=\eta_{k}(\gamma)(v_{1}\wedge\dots\wedge v_{k})$ and let $x:=[v]=\eta_{k}(\gamma)(x_{0})$. By (7) and (8), $x\in\Lambda_{k}(\Gamma)\smallsetminus\operatorname{\mathbb{P}}(\ker T)$. Then by Equation (6), we have that $x=T(x)$, or equivalently, that $T(v)=\lambda v$ for some $\lambda\neq 0$. So
$v$ $\wedge\eta_{d-k}(m)(e_{j_{1}}\wedge\cdots\wedge e_{j_{d-k}})=\frac{1}{\lambda}T(v)\wedge\eta_{d-k}(m)(e_{j_{1}}\wedge\cdots\wedge e_{j_{d-k}})$
$=\frac{1}{\lambda}\det(m)\left(\left(\sum_{1\leq i_{1}<\cdots<i_{k}\leq d}a_{i_{1},\dots,i_{k}}D_{i_{1},\dots,i_{k}}\eta_{k}(\ell)(v)\right)\wedge(e_{j_{1}}\wedge\cdots\wedge e_{j_{d-k}})\right)$
$=\frac{1}{\lambda}\det(m)\left(a_{i_{1}^{\prime},\dots,i_{k}^{\prime}}D_{i_{1}^{\prime},\dots,i_{k}^{\prime}}\eta_{k}(\ell)(v)\right)\wedge(e_{j_{1}}\wedge\cdots\wedge e_{j_{d-k}})=0.$
This contradicts Equation (9). ∎
Lemma 4.9.
Suppose $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a strongly irreducible $\mathsf{P}_{k}$-Anosov subgroup and $\mathsf{G}$ is the real Zariski closure of $\Gamma$. If $\Lambda_{k}(\Gamma)$ is a ${\mathcal{C}}^{1}$-submanifold of $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})$, then $\operatorname{\mathsf{G}}$ has exactly one non-compact simple factor.
Proof.
Let $\operatorname{\mathsf{G}}^{0}$ denote the identity component of $\operatorname{\mathsf{G}}$, and decompose
$\operatorname{\mathsf{G}}^{0}=\mathsf{H}\times\mathsf{L}$
where $\mathsf{H}$ is the product of the non-compact simple factors and $\mathsf{L}$ is the product of the compact simple factors. It suffices to prove that $\mathsf{H}$ is simple.
Let $\Gamma_{0}:=\Gamma\cap\operatorname{\mathsf{G}}^{0}$. By part (2) of Lemma 4.3, there is a non-empty subset $\theta\subset\Delta_{\operatorname{\mathsf{G}}^{0}}$ such that $\Gamma_{0}$ is $\operatorname{\mathsf{P}}_{\theta}$-Anosov, and a smooth embedding $\iota:\operatorname{\mathsf{G}}^{0}/\operatorname{\mathsf{P}}_{\theta}\to\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})$ such that $\iota(\Lambda_{\theta}(\Gamma_{0}))=\Lambda_{k}(\Gamma)$. Since $\Lambda_{k}(\Gamma)$ is a smooth submanifold of $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})$, $\Lambda_{\theta}(\Gamma_{0})$ is a smooth submanifold of $\mathsf{G}^{0}/\operatorname{\mathsf{P}}_{\theta}$, so Theorem 4.6 implies that $\mathsf{P}_{\theta}$ is a maximal parabolic subgroup.
Suppose for a contradiction that $\mathsf{H}=\prod_{j=1}^{m}\mathsf{H}_{j}$ has more than one simple factor. By the maximality of $\operatorname{\mathsf{P}}_{\theta}$, after relabeling the simple factors of $\mathsf{H}$, we can assume that
$\mathsf{P}_{\theta}\cap\mathsf{H}=\mathsf{Q}\times\mathsf{H}_{2}\times\cdots\times\mathsf{H}_{m}\times\mathsf{L},$
where $\mathsf{Q}$ is a parabolic subgroup of $\mathsf{H}_{1}$. It follows that the subgroup $\mathsf{H}_{2}\times\cdots\times\mathsf{H}_{m}\times L$ acts trivially $\mathsf{G}^{0}/\operatorname{\mathsf{P}}_{\theta}$, and hence also acts trivially on $\iota(\Lambda_{\theta}(\Gamma_{0}))=\Lambda_{k}(\Gamma)$. So by Lemma 4.8, $\mathsf{H}_{2}\times\cdots\times\mathsf{H}_{m}\times\mathsf{L}$ is compact which is a contradiction since each $\mathsf{H}_{j}$ is a non-compact simple factor. ∎
5. Rigidity
In this section, we complete the proof of Theorem 1.2 by establishing our rigidity result in the lattice case.
Theorem 5.1.
$\delta^{\alpha_{k}}(\Gamma)=2,$
We originally had a more complicated proof of Theorem 5.1 that was more dynamical in flavor (a sketch of this argument is provided at the end of the section). We thank Andres Sambarino for suggesting the current simpler proof that uses Chow’s theorem [14].
Lemma 5.2.
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (dk)(d-k)-hyperconvex subgroup and $\partial\Gamma$ is a two-sphere, then $\Lambda_{k}(\Gamma)$ is a complex submanifold of $\mathrm{Gr}_{k}(\mathbb{C}^{d})$.
Assuming Lemma 5.2, we will now prove Theorem 5.1.
@@ 5. Rigidity
Let $\mathsf{G}$ denote the real Zariski closure of $\Gamma$ and let $\operatorname{\mathsf{G}}^{0}$ denote the connected component of the identity in $\operatorname{\mathsf{G}}$. Since $\Gamma$ is strongly irreducible, it follows that $\Gamma^{\prime}:=\Gamma\cap\operatorname{\mathsf{G}}^{0}$, and hence also $\operatorname{\mathsf{G}}^{0}$, is irreducible.
Since $\delta^{\alpha_{k}}(\Gamma)=2$, by Corollary 3.2, $\Gamma$ is virtually isomorphic to a uniform lattice in $\mathsf{PSL}(2,\mathbb{C})$, so $\partial\Gamma^{\prime}=\partial\Gamma$ is a two-sphere. We may thus apply Lemma 5.2 to deduce that $\Lambda_{k}(\Gamma)$ is a complex submanifold in $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{C}}^{d})$, which is necessarily bilomorphic to $\mathbb{CP}^{1}$. We thus have an isomorphism of biholomorphism groups
$\mathsf{Aut}_{hol}(\Lambda_{k}(\Gamma))\cong\mathsf{Aut}_{hol}(\mathbb{CP}^{1})=\mathsf{PSL}(2,\operatorname{\mathbb{C}}).$
By Chow’s theorem [14], $\Lambda_{k}(\Gamma)$ is a complex (and also real) algebraic subvariety of $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{C}}^{d})$. Since $\operatorname{\mathsf{G}}$ is the real Zariski closure of $\Gamma$ and $\Lambda_{k}(\Gamma)$ is $\Gamma$-invariant, it follows that $\operatorname{\mathsf{G}}$ preserves $\Lambda_{k}(\Gamma)$. Since $\mathsf{G}\subset\mathsf{PSL}(d,\operatorname{\mathbb{C}})$ acts on $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{C}}^{d})$, and hence on $\Lambda_{k}(\Gamma)$, by biholomorphisms, we have a homomorphism
$q_{k}:\operatorname{\mathsf{G}}\rightarrow\mathsf{Aut}_{hol}(\Lambda_{k}(\Gamma))\cong\mathsf{PSL}(2,\operatorname{\mathbb{C}}).$
By Theorem 4.1, $\operatorname{\mathsf{G}}^{0}$ has no non-trivial proper normal subgroups, which in particular means that $q_{k}|_{\operatorname{\mathsf{G}}^{0}}$ is an isomorphism onto its image. Thus, the Lie algebra $\mathfrak{g}$ of $\mathsf{G}^{0}$ is isomorphic to a sub-algebra of $\mathfrak{sl}(2,\operatorname{\mathbb{C}})$. Also, Theorem 4.1 implies that the Lie algebra $\mathfrak{g}$ is simple. Since $\mathfrak{sl}(2,\operatorname{\mathbb{C}})$ has (real) rank $1$, and has six (real) dimensions, we deduce that either $\mathfrak{g}$ is isomorphic to $\mathfrak{sl}(2,\operatorname{\mathbb{C}})$, or is simple, non-compact, rank $1$, and has dimension at most $5$. By the classification of rank $1$ simple Lie algebras, one deduces that $\mathfrak{g}$ is either isomorphic to $\mathfrak{sl}(2,\operatorname{\mathbb{C}})$ or to $\mathfrak{sl}(2,\operatorname{\mathbb{R}})$.
Since $\operatorname{\mathsf{G}}^{0}$ is connected and isomorphic to a subgroup of $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$, it now follows that $\operatorname{\mathsf{G}}^{0}$ is either isomorphic to $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ or $\mathsf{PSL}(2,\operatorname{\mathbb{R}})$. If $\operatorname{\mathsf{G}}^{0}$ is isomorphic to $\mathsf{PSL}(2,\operatorname{\mathbb{R}})$, then by the Milnor-Svarc lemma, $\partial\Gamma^{\prime}$ is a subset of $\mathbb{S}^{1}$, which contradicts the fact that $\partial\Gamma^{\prime}=\partial\Gamma$ is a $2$-sphere. Hence, we must have that $\operatorname{\mathsf{G}}^{0}$ is isomorphic to $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$.

### v2

@@ 4. Rigidity
Let $\mathsf{G}\subset\mathsf{PSL}(d,\operatorname{\mathbb{C}})$ denote the real Zariski closure of $\Gamma$ and let $\operatorname{\mathsf{G}}^{0}$ denote the connected component of the identity in $\operatorname{\mathsf{G}}$. Since $\Gamma$ is strongly irreducible, it follows that $\Gamma^{\prime}:=\Gamma\cap\operatorname{\mathsf{G}}^{0}$, and hence also $\operatorname{\mathsf{G}}^{0}$, is irreducible.
Since $\delta^{\alpha_{k}}(\Gamma)=2$, by Corollary 3.2, $\Gamma$ is virtually isomorphic to a uniform lattice in $\mathsf{PSL}(2,\mathbb{C})$, so $\partial\Gamma^{\prime}=\partial\Gamma$ is a two-sphere. We may thus apply Lemma 4.2 to deduce that $\Lambda_{k-1,k,k+1}(\Gamma)$ is a complex submanifold in $\operatorname{\mathcal{F}}_{k-1,k,k+1}(\operatorname{\mathbb{C}}^{d})$, which is necessarily biholomorphic to $\mathbb{CP}^{1}$. We thus have an isomorphism of biholomorphism groups
$\mathsf{Aut}_{hol}(\Lambda_{k-1,k,k+1}(\Gamma))\cong\mathsf{Aut}_{hol}(\mathbb{CP}^{1})=\mathsf{PSL}(2,\operatorname{\mathbb{C}}).$
In what follows we identify
$\mathsf{Aut}_{hol}(\Lambda_{k-1,k,k+1}(\Gamma))=\mathsf{PSL}(2,\operatorname{\mathbb{C}}).$
By Chow’s theorem [13], $\Lambda_{k-1,k,k+1}(\Gamma)$ is a complex (and also real) algebraic subvariety of $\operatorname{\mathcal{F}}_{k-1,k,k+1}(\operatorname{\mathbb{C}}^{d})$. Since $\operatorname{\mathsf{G}}$ is the real Zariski closure of $\Gamma$ and $\Lambda_{k-1,k,k+1}(\Gamma)$ is $\Gamma$-invariant, it follows that $\operatorname{\mathsf{G}}$ preserves $\Lambda_{k-1,k,k+1}(\Gamma)$. Since $\mathsf{G}\subset\mathsf{PSL}(d,\operatorname{\mathbb{C}})$ acts on $\operatorname{\mathcal{F}}_{k-1,k,k+1}(\operatorname{\mathbb{C}}^{d})$, and hence on $\Lambda_{k-1,k,k+1}(\Gamma)$, by biholomorphisms, this $\operatorname{\mathsf{G}}$-action defines a homomorphism
$q_{k}:\operatorname{\mathsf{G}}\rightarrow\mathsf{Aut}_{hol}(\Lambda_{k-1,k,k+1}(\Gamma))=\mathsf{PSL}(2,\operatorname{\mathbb{C}}).$
Claim: $q_{k}(\operatorname{\mathsf{G}}^{0})=\mathsf{PSL}(2,\operatorname{\mathbb{C}})$.
@@
Since $\operatorname{\mathsf{G}}^{0}$ is irreducible, we observed earlier that it is semisimple and its center is trivial, so we can decompose it into a product of simple Lie groups. Fix a simple factor $\operatorname{\mathsf{G}}_{1}$ such that $q_{k}(\operatorname{\mathsf{G}}_{1})$ is non-trivial. Then $q_{k}(\operatorname{\mathsf{G}}_{1})$ is normal in $q_{k}(\operatorname{\mathsf{G}}^{0})=\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ and $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ is simple, so $q_{k}(\operatorname{\mathsf{G}}_{1})=\mathsf{PSL}(2,\operatorname{\mathbb{C}})$. Further, since $\operatorname{\mathsf{G}}_{1}$ is simple and has trivial center (since $\operatorname{\mathsf{G}}^{0}$ has trivial center), we see that $q_{k}|_{\operatorname{\mathsf{G}}_{1}}:\operatorname{\mathsf{G}}_{1}\to\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ is an isomorphism. Since $q_{k}(\operatorname{\mathsf{G}}_{1})=\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ commutes with the images of the other simple factors and $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ has trivial center, $\ker q_{k}$ coincides with the product of the other simple factors. Hence we can write $\operatorname{\mathsf{G}}^{0}$ as the direct product
$\operatorname{\mathsf{G}}^{0}=\operatorname{\mathsf{G}}_{1}\times\operatorname{\mathsf{G}}_{2}$
where $\operatorname{\mathsf{G}}_{2}=\ker q_{k}$. We will show that $\operatorname{\mathsf{G}}_{2}=1$, which will imply that $q_{k}:\operatorname{\mathsf{G}}^{0}\rightarrow\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ is an isomorphism.
For $j=1,2$, let $\operatorname{\mathsf{G}}_{j}^{\prime}$ denote the preimage of $\operatorname{\mathsf{G}}_{j}$ under the quotient map $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{C}})\rightarrow\mathsf{PSL}(d,\operatorname{\mathbb{C}})$. Notice that $\operatorname{\mathsf{G}}_{1}^{\prime}$ and $\operatorname{\mathsf{G}}_{2}^{\prime}$ are commuting subgroups of $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{C}})$, so by Lemma 4.3, there exist integers $d_{1}$ and $d_{2}$, representations $\rho_{1}:\operatorname{\mathsf{G}}_{1}^{\prime}\rightarrow\operatorname{\mathsf{SL}}(d_{1},\operatorname{\mathbb{C}})$ and $\rho_{2}:\operatorname{\mathsf{G}}_{2}^{\prime}\rightarrow\operatorname{\mathsf{SL}}(d_{2},\operatorname{\mathbb{C}})$ and an identification
$\operatorname{\mathbb{C}}^{d}=\operatorname{\mathbb{C}}^{d_{1}}\otimes\operatorname{\mathbb{C}}^{d_{2}}$
such that an element $g=g_{1}g_{2}\in\operatorname{\mathsf{G}}_{1}^{\prime}\operatorname{\mathsf{G}}_{2}^{\prime}$ acts by
$g(v\otimes w)=(\rho_{1}(g_{1})(v))\otimes(\rho_{2}(g_{2})(w)).$
Claim: If $\ell\in\{k-1,k,k+1\}\cap\{1,\dots,d-1\}$, then
$\Lambda_{\ell}(\Gamma)\subset\{V\otimes\operatorname{\mathbb{C}}^{d_{2}}:V\in\operatorname{\mathrm{Gr}}_{\ell/d_{2}}(\operatorname{\mathbb{C}}^{d_{1}})\}.$
Proof of Claim: Let $\gamma\in\Gamma^{\prime}$ be an infinite order element and let $U_{\gamma}^{+}\in\Lambda_{\ell}(\Gamma)$ be the attracting fixed point of $\gamma$. Then $U_{\gamma}^{+}$ is a direct sum of generalized eigenspaces of $\gamma$ (corresponding to the largest $\ell$ eigenvalues). If we write $\gamma=\gamma_{1}\gamma_{2}$ where $\gamma_{j}\in\operatorname{\mathsf{G}}_{j}$, then
$U_{\gamma}^{+}=\oplus_{i}V_{i}\otimes W_{i}$
where each $V_{i}$ is a generalized eigenspace of $\rho_{1}(\gamma_{1})$ and each $W_{i}$ is a generalized eigenspace of $\rho_{2}(\gamma_{2})$. Since $\operatorname{\mathsf{G}}_{2}=\ker q_{k}$, for all $g_{2}\in\operatorname{\mathsf{G}}_{2}$, we have $g_{2}(U_{\gamma}^{+})=U_{\gamma}^{+}$, so the union of subspaces $\cup_{i}W_{i}$ is invariant under $\rho_{2}(g_{2})$. Since $\rho_{2}$ is strongly irreducible $W_{i}=\operatorname{\mathbb{C}}^{d_{2}}$ for all $i$. Then
$U_{\gamma}^{+}=V\otimes\operatorname{\mathbb{C}}^{d_{2}}$
[…]
The above claim implies that there is some $\ell\in\{k-1,k\}$ such that $\frac{\ell}{d_{2}}$ and $\frac{\ell+1}{d_{2}}$ are both integers, so we must have $d_{2}=1$. This in turn implies that $\operatorname{\mathsf{G}}_{2}$ is a finite, normal subgroup of $\operatorname{\mathsf{G}}^{0}$. Since $\operatorname{\mathsf{G}}^{0}$ is connected, its action by conjugation on $\operatorname{\mathsf{G}}_{2}$ is necessarily trivial, so $\operatorname{\mathsf{G}}_{2}$ must be in the center of $\operatorname{\mathsf{G}}^{0}$, which is trivial. Thus, $\operatorname{\mathsf{G}}^{0}=\operatorname{\mathsf{G}}_{1}$ and $q_{k}:\operatorname{\mathsf{G}}^{0}\rightarrow\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ is an isomorphism.
To finish the proof, we need to show that $\operatorname{\mathsf{G}}=\operatorname{\mathsf{G}}^{0}$. Since $\operatorname{\mathsf{G}}^{0}\subset\mathsf{PSL}(d,\operatorname{\mathbb{C}})$ is irreducible and isomorphic to $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$, we may view it as the image of an injective irreducible representation $\tau:\mathsf{PSL}(2,\operatorname{\mathbb{C}})\to\mathsf{PSL}(d,\operatorname{\mathbb{C}})$. By the classification of irreducible representations of $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$, there is a linear identification between $\operatorname{\mathbb{C}}^{d}$ and the symmetric tensor ${\rm Sym}^{d-1}(\operatorname{\mathbb{C}}^{2})$ with the defining property that